Solve
dy/dx + xy=y^2×e^(x^2/2)×cosx
1
Expert's answer
2020-11-02T20:05:17-0500

"y' + xy = {y^2}{e^{\frac{{{x^2}}}{2}}}\cos x"

Let

"y = uv \Rightarrow y' = u'v + uv'"

Then

"\begin{array}{l}
u'v + uv' + xuv = {u^2}{v^2}{e^{\frac{{{x^2}}}{2}}}\cos x\\
u'v + u\left( {v' + xv} \right) = {u^2}{v^2}{e^{\frac{{{x^2}}}{2}}}\cos x
\end{array}"

Let

"v' + xv = 0 \Rightarrow \frac{{dv}}{{dx}} = - xv \Rightarrow \frac{{dv}}{v} = - xdx \Rightarrow \ln v = - \frac{{{x^2}}}{2} \Rightarrow v = {e^{ - \frac{{{x^2}}}{2}}}"

Then

"u'{e^{ - \frac{{{x^2}}}{2}}} = {u^2}{e^{ - {x^2}}}{e^{\frac{{{x^2}}}{2}}}\cos x"

"\begin{array}{l}
u' = {u^2}{e^{ - {x^2}}}{e^{{x^2}}}\cos x\\
u' = {u^2}\cos x
\end{array}"

"\begin{array}{l}
\frac{{du}}{{dx}} = {u^2}\cos x\\
\frac{{du}}{{{u^2}}} = \cos xdx\\
 - \frac{1}{u} = \sin x - C\\
u = \frac{1}{{C - \sin x}}
\end{array}"

"y = uv = \frac{{{e^{ - \frac{{{x^2}}}{2}}}}}{{C - \sin x}}"

Answer: "y = \frac{{{e^{ - \frac{{{x^2}}}{2}}}}}{{C - \sin x}}"